G-İBrd, Gee rare, and Ss In order, rou a die and the ame is endede oht 1s the prni Lu1w
Answer #1

For anyone in any particular turn:

P(Winning) = P(Rolling an even number) = 3/6 = 1/2

So P(Not winning) = 1 - 1/2 = 1/2

Bill can win in 1st, 4th, 7th,..... rounds because his turn is coming after every three turn.

Hence,

P(Bill wins)

=(1/2) + (1/2)^4 + (1/2)^7 + ....

=rac{(1/2)}{1 - (1/2)^3}

(7/8)

= 4/7